# Euler systems and relative Satake isomorphism

### v1

@@ 1.3. Related works
In our language, the gap in Cornut’s construction [Cor18] (see [Sha23]) is a difference between functions on $\mathbf{X}$ and distributions on $\mathbf{X}$. In this paper, we establish tame norm relations at split primes. We believe our framework also works in the original non-split setting studied by Cornut and we plan to pursue such a generalization in a future paper (see §1.4.1 below).
In the other works cited in Table 1, the Euler system is constructed using the zeta integral method first developed by Loeffler–Skinner–Zerbes [LSZ22b]. In this method, one writes down a carefully chosen candidate for the test vector in $C_{c}^{\infty}(X,\mathbb{Z}_{p})^{K^{1}}$ and verifies that its trace is equal to $\mathcal{H}_{\ell}\cdot\Phi_{0}$ by an explicit, often intricate, zeta integral computation. It also relies crucially on local multiplicity one. Our method avoids both the computations and the local multiplicity one hypothesis.
In certain spherical settings where classes are obtained by a pushforward construction, Loeffler gave a construction of wild norm relations [Loe21]. He also considered more general cases where a mirabolic subgroup has an open orbit (Definition 4.1.1 of op. cit.). In many such cases, we reinterpret the mirabolic subgroup as the point stabilizer of a naturally occurring affine inhomogeneous spherical variety. However, in some examples such as $(\mathrm{GSp}_{4},\mathrm{GL}_{2}\times_{\mathbb{G}_{m}}\mathrm{GL}_{2})$, only a parabolic subgroup of $\mathbf{G}$ has an open orbit. Our method does not handle this case. This is an Eisenstein degeneration of the spherical pair $(\mathrm{SO}_{4}\times\mathrm{SO}_{5},\mathrm{SO}_{4})$, in a precise sense explained in [LR24], so it would be interesting to understand how this can be interpreted as an operation on spherical varieties.
@@ 1.4.1. Removing conditions
It would also be useful to have a non-split version of our results. Indeed, in certain twisted settings, even the construction of a split anticyclotomic Euler system requires non-split groups. The necessary local harmonic analysis results should be within reach of current methods. For example, [SW22] only assumes quasi-split, and [CZ23] handles many symmetric cases.
1.4.2. Speculations
One byproduct of our observation of automatic divisibility (Proposition 3.9) is that there should be some modification to the local unramified conjecture [BZSV24, Conjecture 7.5.1] when the coefficient ring is not a field, cf. §9.4 of op. cit. We hope this provides a useful piece of phenomenon in the future refinement of this framework.
@@ 1.4.2. Speculations
in larger generality. Indeed, replacing “integral motivic classes” with “automorphic functions”, then this space (though not its two constituent pieces) has a conjectural dual description in the relative Langlands program. In addition to the pushforward of cycles or motivic classes described above, the arithmetic theta lifting (cf. [Liu11, LL21, Dis24]) should also be part of this framework. It would also be interesting to understand its relation with the recent boundary class construction of Skinner–Sangiovanni-Vincentelli [SV24].
@@ Acknowledgments
It is clear that most of this paper is built on the fundamental works of Yiannis Sakellaridis, and we would like to thank him in particular for some very insightful conversations.
We would also like to thank David Ben-Zvi, Ashay Burungale, Christopher Skinner, Ye Tian and Wei Zhang for their interests and helpful discussions.
The first and second named authors are supported by the National Key R&\&&D Program of China No. 2023YFA1009702 and the National Natural Science Foundation of China No. 12371012. The second named author is also supported by National Natural Science Foundation of Beijing, China No. 24A10020.
@@ 2. Spherical varieties
Let $\mathbf{G}$ be a split reductive group with a Borel subgroup $\mathbf{B}$. Let $\mathbf{X}$ be a variety with a right $\mathbf{G}$-action. Recall that $\mathbf{X}$ is spherical if $\mathbf{B}\subseteq\mathbf{G}$ acting on $\mathbf{X}$ has an open orbit. We make the following assumptions
@@
Assumption (4) is included also for simplicity of notation. It will hold if $\mathbf{H}$ is reductive or if $\mathbf{X}$ is of the form $\mathbf{G}\times^{\mathbf{H}^{\prime}}V$, where $\mathbf{H}^{\prime}$ is reductive and $V$ is a linear representation of $\mathbf{H}^{\prime}$. These are the two cases needed for the Euler system constructions in the paper. In general, one may assume there is an $\mathbf{G}$-eigenmeasure after a trivial modification [Sak08, §3.8], and our formulae in §3.3 need to be modified by the corresponding character in a well-understood way.
@@ 2.1.1. Notations
Write $\Lambda_{X}$ for its cocharacter lattice, and let $\mathfrak{a}_{X}=\Lambda_{X}\otimes_{\mathbb{Z}}\mathbb{Q}$. Similarly define $\mathfrak{a}$, then we have a quotient map $\mathfrak{a}\twoheadrightarrow\mathfrak{a}_{X}$. There is a natural cone $\mathcal{V}\subseteq\mathfrak{a}_{X}$ containing the image of the negative Weyl chamber. Let $\Lambda_{X}^{+}=\mathcal{V}\cap\Lambda_{X}$. Its elements will be called $\mathbf{X}$-anti-dominant.
@@
Knop and Schalke defined a dual group $\check{G}_{X}$ whose coroot system is $(X^{*}(\mathbf{A}_{X}),\Phi_{X})$ [KS17]. It is a subgroup of the Langlands dual group $\check{G}$ of $\mathbf{G}$. We have an equality $\check{G}_{X}=\check{G}$ only if for all simple roots $\alpha$, the pair $(\mathring{\mathbf{X}},\alpha)$ is of type T in the classification below. We call $\mathbf{X}$ with this property strongly tempered.
@@ 2.1.2. Classification of roots
  • Type G: $\ast=\mathrm{PGL}_{2}\backslash\mathrm{PGL}_{2}$.
  • Type U: $\mathbf{S}\backslash\mathrm{PGL}_{2}$, where $\mathbf{S}$ is the subgroup of a Borel subgroup which contains the unipotent radical.
  • Type T: $\mathbf{T}\backslash\mathrm{PGL}_{2}$, where $\mathbf{T}$ is a maximal torus.
  • Type N: $\mathbf{N}(\mathbf{T})\backslash\mathrm{PGL}_{2}$, where $\mathbf{N}(\mathbf{T})$ is the normalizer of the torus. This case is excluded by our assumption.
@@ 2.2. Equivariant bundles
For our intended application, it is necessary to consider certain torus bundles over $\mathbf{X}$.
@@
It is combinatorially trivial if and only if $\nu|_{\mathbf{H}\cap\mathbf{A}}=1$, where recall that $\mathbf{A}$ is a maximal torus in the Borel subgroup $\mathbf{B}$ such that $\mathbf{H}\mathbf{B}$ is open in $\mathbf{G}$. This is the condition imposed by Loeffler for wild norm relations [Loe21, §4.6].
@@
On the representation theory side, this means if $\pi$ is $\mathbf{X}$-distinguished, then $\pi\times\chi$ is $\widetilde{\mathbf{X}}$-distinguished for any character $\chi$ of the torus. So we are looking a family of character twists for $\pi$, exactly what is needed for Euler system constructions. If $\mathbf{T}=\mathbb{G}_{m}$, this is related to having an “$s$-variable” in the $L$-function.
The existence of such a bundle is unfortunately quite restrictive, and it rules out very interesting cases, including the triple product case $\mathbf{G}=\mathrm{SO}_{3}\times\mathrm{SO}_{4}$, $\mathbf{H}=\mathrm{SO}_{3}$.
@@
Let $\mathbf{Y}$ be a $\mathbf{B}$-orbit in $\mathbf{X}$ of maximal rank. Let $\alpha$ be a simple root of $\mathbf{G}$ with associated parabolic $\mathbf{P}_{\alpha}$. The pair $(\mathbf{Y},\alpha)$ cannot have type T, since otherwise there would be a spherical root of type T. It follows that the pair is of type G or U. In particular, $\mathbf{Y}\mathbf{P}_{\alpha}$ is either $\mathbf{Y}$ or the disjoint union of an open $\mathbf{B}$-orbit and a closed $\mathbf{B}$-orbit. Moreover, all $\mathbf{B}$-orbits are of maximal rank, so all pairs $(\mathbf{Y},\alpha)$ are not of type T.
@@
Since $\mathbf{P}_{\alpha}\subseteq\mathbf{B}$, the restriction of $\nu$ to the first term is trivial, so it descends to a character $\bar{\nu}$ on the quotient. On the other hand, this quotient is the stabilizer group in the $\mathrm{PGL}_{2}$-spherical variety $\mathbf{Y}\mathbf{P}_{\alpha}/\mathcal{R}(\mathbf{P}_{\alpha})$, which is of type U. But then $\bar{\nu}$ is trivial by Example 2.3. Therefore, $\nu|_{\mathbf{H}\cap\mathbf{P}_{\alpha}}=1$. In other words, $\widetilde{\mathbf{X}}$ trivializes as a $\mathbf{B}_{\alpha}$-bundle over $\mathbf{Y}\mathbf{P}_{\alpha}$.
@@ 3. Local computations
The group $G$ has a natural smooth left action on the function space $C^{\infty}(X,\mathbb{C})$. In this section, we recall its spectral decomposition, following the works of Sakellaridis and his collaborators. The goal is to observe certain automatic divisibility properties and match them with a corresponding phenomenon in geometry.
@@ 3.1. Assumptions
For all results in this section, we need to impose some “good reduction” hypotheses, which we will specify. In the global setting, they hold for all but finitely many places.
@@ 3.2.2. Interaction with bundle
Clearly, $K^{1}$ is a normal subgroup of $K^{0}$ whose quotient is $\mathbb{F}^{\times}$, which has size $q-1$. The main result of this subsection is to describe the space $\widetilde{X}/K^{1}$.
@@
Let $x\in{\widetilde{X}}$. If the action of $K^{1}$ on $xK^{0}$ has fewer than $q-1$ orbits, then $x$ lies on a wall of type T.
@@
In this proof, we will only work over the generic fibre, so the standard theory of spherical embeddings apply without comment. Further note that we are only working over the open $\mathbf{G}$-orbit $\mathbf{X}^{\bullet}$, so we will assume that $\mathbf{X}$ is homogeneous.
Over the open orbit $\mathring{\mathbf{X}}$, fix a trivialization of $\widetilde{\mathbf{X}}$. Identify $A_{X}$ with a subset of $\mathring{X}$ using the orbit map. Let $x\in A_{X}^{+}$ be an element such that $K^{1}$ splits $xK^{0}$ into less than $q-1$ pieces. Then there exists $g\in K$ such that $(x,1)g=(x,\alpha)$, where $\alpha\not\equiv 1\pmod{\varpi}$. For any $n\geq 1$, we have
@@
The combinatorially trivial $\mathbb{G}_{m}$-bundle $\widetilde{\mathbf{X}}$ also extends to the boundary using the compactification given by the same fan decomposition. By [SV17, Proposition 2.3.8(3)], the spherical roots of $\mathbf{Z}$ are exactly $\Theta$, so $\mathbf{Z}$ is a spherical variety with only spherical roots of type G. Proposition 2.5 implies that the bundle over $\mathbf{Z}$ must split. On the other hand, taking limit as $n\to\infty$ in equation (2), we see that $g$ translates between the $q-1$ disks above the point $x^{\infty}$. This is a contradiction, proving that $\check{\lambda}$ must lie on a wall of type T. ∎
@@ 3.3. Relative Satake isomorphism
In this subsection, we will introduce the relative Satake isomorphism. In the homogeneous case, the results are unconditional and due to Sakellaridis [Sak13, Sak18]. In a second case, which we call “totally type T”, the analogous result holds in the equal-characteristic case [SW22], and we conjecture that it carries over to the mixed characteristic case.
@@ 3.3.1. Hecke algebra and Satake isomorphism
Fix the Haar measure on $G$ so that $K$ has volume 1, then we can define the Hecke algebra $\mathcal{H}(G):=C_{c}^{\infty}(K\backslash G/K,\mathbb{C})$ whose algebra structure is given by convolution
@@
[…]
Note that when $X=G$, this action is different from the convolution.
@@
[…]
where $\Lambda=X_{*}(\mathbf{A})=\mathbf{A}(F)/\mathbf{A}(\mathcal{O})$. By restriction, we have a natural ring homomorphism $\mathbb{C}[\Lambda]^{W}\to\mathbb{C}[\Lambda_{X}^{+}]$. Let $A^{*}$ be the set of unramified characters of $A$, so $A^{*}=X^{*}(\mathbf{A})\otimes_{\mathbb{Z}}\mathbb{C}^{\times}$ is an algebraic torus. There is a Fourier transform

### v2

@@ 1.5. Related works
1.5.1. Zeta integral approach
In all works cited in Table 1 (except [LLZ14]), the tame norm relation is verified using the zeta integral method first developed by Loeffler–Skinner–Zerbes [LSZ22b]. In this method, one writes down a carefully chosen candidate for the test vector in $C_{c}^{\infty}(X,\mathbb{Z}_{p})^{K^{1}}$ and verifies that its trace is equal to $\mathcal{H}_{\ell}\cdot\Phi_{0}$ by an explicit, often intricate, zeta integral computation. It also relies crucially on local multiplicity one. Our method avoids both the computations and the local multiplicity one hypothesis.
1.5.2. Other approaches
The idea of using spherical varieties in the construction of Euler systems goes back to Cornut [Cor18]. Unfortunately, there is a gap in his construction related to the definition of Hecke operators, as explained in [Sha23]. From the point of view of our framework, his setting presents additional features worthy of further investigation (cf. Remark 3.19).
In certain spherical settings where classes are obtained by a pushforward construction, Loeffler gave a systematic construction of wild norm relations [Loe21]. He also considered more general cases where a mirabolic subgroup has an open orbit (Definition 4.1.1 of op. cit.). In many such cases, we reinterpret the mirabolic subgroup as the point stabilizer of a naturally occurring affine inhomogeneous spherical variety. However, in some examples such as $(\mathrm{GSp}_{4},\mathrm{GL}_{2}\times_{\mathbb{G}_{m}}\mathrm{GL}_{2})$, only a parabolic subgroup of $\mathbf{G}$ has an open orbit. Our method does not handle this case. This is an Eisenstein degeneration of the spherical pair $(\mathrm{SO}_{4}\times\mathrm{SO}_{5},\mathrm{SO}_{4})$, and it would be interesting to understand how this can be interpreted as an operation on spherical varieties.
@@ 1.5.3. Removing conditions
It would also be useful to more systematically develop our results in the non-split setting. The necessary local harmonic analysis results should be within reach of current methods. For example, [SW22] only assumes that the group is quasi-split, and [CZ23] handles many symmetric cases.
1.5.4. Vista
@@
in larger generality. Indeed, replacing “integral motivic classes” with “automorphic functions”, then this space (though not its two constituent pieces) has a conjectural dual description in the relative Langlands program. In addition to the pushforward of cycles or motivic classes described above, the arithmetic theta lifting (cf. [Liu11, LL21, Dis24]) should also be part of this framework. It would also be interesting to understand its relation with the recent relative cohomology construction of Sangiovanni–Skinner, cf. [SVS24].
@@ Acknowledgments
It is clear that most of this paper is built on the fundamental works of Yiannis Sakellaridis. We would like to thank him for insightful conversations and for pointing out a major gap in an earlier version of the paper.
We would also like to thank David Ben-Zvi, Ashay Burungale, David Loeffler, Christopher Skinner, Ye Tian and Wei Zhang for their interests and helpful discussions.
The first and second named authors are supported by the National Key R&\&D Program of China No. 2023YFA1009702 and the National Natural Science Foundation of China No. 12371012. The second named author is also supported by National Natural Science Foundation of Beijing, China No. 24A10020.
@@ 2. Spherical varieties
Let $\mathbf{G}$ be a split reductive group with a Borel subgroup $\mathbf{B}$. Let $\mathbf{X}$ be a variety with a right $\mathbf{G}$-action. Recall that $\mathbf{X}$ is spherical if the action of $\mathbf{B}\subseteq\mathbf{G}$ on $\mathbf{X}$ has an open orbit. We make the following assumptions
@@
The condition that $\mathbf{G}$ is split is present in most of the literature on spherical varieties. Even though our groups are non-split globally, we will only apply the results of this section when $F$ is a local field where the group $\mathbf{G}$ splits
@@
Assumption (4) is also included only for simplicity of notation. It holds if $\mathbf{H}$ is reductive or if $\mathbf{X}$ is of the form $\mathbf{G}\times^{\mathbf{H}^{\prime}}V$, where $\mathbf{H}^{\prime}$ is reductive and $V$ is a linear representation of $\mathbf{H}^{\prime}$. These are the two cases needed for the Euler system constructions in the paper. In general, one may assume there is an $\mathbf{G}$-eigenmeasure after trivially modifying [Sak08, §3.8], and our formulae in §3.3 need to be modified by the corresponding character in a well-understood way.
@@ 2.1.1. Notations
Write $\Lambda_{X}$ for its cocharacter lattice, and let $\mathfrak{a}_{X}=\Lambda_{X}\otimes_{\mathbb{Z}}\mathbb{Q}$. Similarly define $\mathfrak{a}$ as the coroot lattice of $\mathbf{G}$ tensored with $\mathbb{Q}$, then we have a quotient map $\mathfrak{a}\twoheadrightarrow\mathfrak{a}_{X}$. There is a natural cone $\mathcal{V}\subseteq\mathfrak{a}_{X}$ containing the image of the negative Weyl chamber. Let $\Lambda_{X}^{+}=\mathcal{V}\cap\Lambda_{X}$. Its elements will be called $\mathbf{X}$-anti-dominant.
@@
Knop and Schalke defined a dual group $\check{G}_{X}$ whose coroot system is $(X^{*}(\mathbf{A}_{X}),\Phi_{X})$ [KS17]. It is a subgroup of the Langlands dual group $\check{G}$ of $\mathbf{G}$. We have an equality $\check{G}_{X}=\check{G}$ only if for all simple roots $\alpha$, the pair $(\mathring{\mathbf{X}},\alpha)$ is of type T in the classification below. Spherical varieties $\mathbf{X}$ with this property are said to be strongly tempered.
@@ 2.1.2. Classification of roots
[…]
@@ 2.2. Equivariant bundles
For applications to Euler systems, it is necessary to consider certain torus bundles over $\mathbf{X}$.
@@
It is combinatorially trivial if and only if $\nu|_{\mathbf{H}\cap\mathbf{A}}=1$, where recall that $\mathbf{A}$ is a maximal torus in the Borel subgroup $\mathbf{B}$ such that $\mathbf{H}\mathbf{B}$ is open in $\mathbf{G}$. This is the condition imposed by Loeffler when he systematically studied wild norm relations in [Loe21, §4.6].
@@
On the representation theory side, this means if $\pi$ is $\mathbf{X}$-distinguished, then $\pi\times\chi$ is $\widetilde{\mathbf{X}}-$distinguished for any character $\chi$ of the torus. So we are looking a family of character twists for $\pi$, exactly what is needed for Euler system constructions. If $\mathbf{T}=\mathbb{G}_{m}$, then this is related to having an “$s$-variable” in the $L$-function.
The existence of such a bundle is unfortunately quite restrictive, and it rules out very interesting cases, including the triple product case $\mathbf{G}=\mathrm{SO}_{3}\times\mathrm{SO}_{4}$, $\mathbf{H}=\mathrm{SO}_{3}$. Indeed, the Ichino formula only calculates the central value, and $\mathbf{H}$ does not admit non-trivial characters in this case.
The following technical proposition plays a crucial role in understanding the divisibility properties needed to characterize the trace map considered in the next section. Its proof consists of expanding the $\mathbf{B}$-trivialization over the open orbit to the entire variety using simple reflections.
@@
Let $\mathbf{Y}$ be a $\mathbf{B}$-orbit in $\mathbf{X}$ of maximal rank. Let $\alpha$ be a simple root of $\mathbf{G}$ with associated parabolic $\mathbf{P}_{\alpha}$. The pair $(\mathbf{Y},\alpha)$ cannot have type T, since otherwise there would be a spherical root of type T. It follows that the pair is of type G or U. In particular, $\mathbf{Y}\mathbf{P}_{\alpha}$ is either $\mathbf{Y}$ or the disjoint union of an open $\mathbf{B}$-orbit and a closed $\mathbf{B}$-orbit. In the latter case, both $\mathbf{B}$-orbits have the same rank. It follows that all $\mathbf{B}$-orbits are of maximal rank, so all pairs $(\mathbf{Y},\alpha)$ are not of type T.
@@
Since $\mathcal{R}(\mathbf{P}_{\alpha})\subseteq\mathbf{B}$, the restriction of $\nu$ to the first term is trivial, so it descends to a character $\bar{\nu}$ on the quotient. On the other hand, this quotient is the stabilizer group in the $\mathrm{PGL}_{2}$-spherical variety $\mathbf{Y}\mathbf{P}_{\alpha}/\mathcal{R}(\mathbf{P}_{\alpha})$, which is necessarily of type U. But then $\bar{\nu}$ is trivial by Example 2.3. Therefore, $\nu|_{\mathbf{H}\cap\mathbf{P}_{\alpha}}=1$. In other words, $\widetilde{\mathbf{X}}$ trivializes as a $\mathbf{B}_{\alpha}$-bundle over $\mathbf{Y}\mathbf{P}_{\alpha}$.
@@ 3. Local computations
The group $G$ has a natural smooth left action on the function space $C^{\infty}(X,\mathbb{C})$. In this section, we recall the spectral decomposition of its unramified part, following the works of Sakellaridis and his collaborators. The goal is to observe certain automatic divisibility properties and match them with a corresponding phenomenon in geometry.
@@ 3.1. Assumptions
For all results in this section, we need to impose the following “good reduction” hypotheses. In the global setting, they hold for all but finitely many places.
@@ 3.2.2. Interaction with bundle
Clearly, $K^{1}$ is a normal subgroup of $K^{0}$ whose quotient is $\mathbb{F}^{\times}$, which has size q1q-1. The main result of this subsection is to describe the space $\widetilde{X}/K^{1}$.
@@
We are interested in understanding the image of the trace map
$\operatorname{Tr}_{K_{0}}^{K^{1}}:C_{c}^{\infty}(\widetilde{X},\mathbb{Z}_{p})^{K^{1}}\to C_{c}^{\infty}(\widetilde{X},\mathbb{Z}_{p})^{K^{0}},$
which is equivalent to studying when restricting to $K^{1}$ splits an $K^{0}$-orbit in the generalized Cartan decomposition. We give a partial implication, which we suspect can be upgraded to an equivalence. The proof is based on considering asymptotic degenerations of the bundle $\widetilde{\mathbf{X}}\to\mathbf{X}$.
@@
Let $x\in{\widetilde{X}}$. If the action of $K^{1}$ on $xK^{0}$ has fewer than q1q-1 orbits, then $x$ lies on a wall of type T.
@@
In this proof, we will only work over the generic fibre, so the standard theory of spherical embeddings applies without comment. Further note that we are only working over the open $\mathbf{G}$-orbit $\mathbf{X}^{\bullet}$, so we will assume that $\mathbf{X}$ is homogeneous.
Over the open orbit $\mathring{\mathbf{X}}$, fix a trivialization
$\widetilde{\mathbf{X}}\times_{\mathbf{X}}\mathring{\mathbf{X}}\simeq\mathring{\mathbf{X}}\times\mathbb{G}_{m}$
as a $\mathbf{B}$-equivariant bundle. Identify $A_{X}$ with a subset of $\mathring{X}$ using the orbit map. Let $x\in A_{X}^{+}$ be an element such that $K^{1}$ splits $xK^{0}$ into less than q1q-1 pieces. Then there exists $g\in K$ such that $(x,1)g=(x,\alpha)$, where $\alpha\not\equiv 1\pmod{\varpi}$. For any $n\geq 1$, we have
@@
The combinatorially trivial $\mathbb{G}_{m}$-bundle $\widetilde{\mathbf{X}}$ also extends to the boundary using the compactification given by the same fan decomposition. By [SV17, Proposition 2.3.8(3)], the spherical roots of $\mathbf{Z}$ are exactly $\Theta$, so $\mathbf{Z}$ is a spherical variety with only spherical roots of type G. Proposition 2.5 implies that the bundle over $\mathbf{Z}$ must split. On the other hand, taking limit as $n\to\infty$ in equation (2), we see that $g$ translates between the q1q-1 disks above the point $x^{\infty}$. This is a contradiction, proving that $\check{\lambda}$ must lie on a wall of type T. ∎
@@ 3.3. Relative Satake isomorphism
In this subsection, we will introduce the relative Satake isomorphism. In the homogeneous case, the results are unconditional and due to Sakellaridis [Sak13, Sak18]. In the strongly tempered case, the analogous result holds in the equal-characteristic case [SW22], and we conjecture that it carries over to the mixed characteristic case.
@@ 3.3.1. Hecke algebra and Satake isomorphism
$(f\cdot\phi)(x)=\int_{G}\phi(xg)f(g)\,dg.$
Note that when $X=G$, this action is different from the convolution by an inverse.
@@
(3) $\mathcal{H}(G)\xrightarrow{\,\sim\,}\mathcal{H}(A)^{W}\simeq\mathbb{C}[\Lambda]^{W},\quad f\mapsto\hat{f}(t):=\delta(t)^{\frac{1}{2}}\int_{N}f(tn)dn$
where $\Lambda=X_{*}(\mathbf{A})=\mathbf{A}(F)/\mathbf{A}(\mathcal{O})$, and $\delta$ is the modulus character of the Borel subgroup containing $N$. By restriction, we have a natural ring homomorphism $\mathbb{C}[\Lambda]^{W}\to\mathbb{C}[\Lambda_{X}^{+}]$. Let $A^{*}$ be the set of unramified characters of $A$, so $A^{*}=X^{*}(\mathbf{A})\otimes_{\mathbb{Z}}\mathbb{C}^{\times}$ is an algebraic torus. There is a Fourier transform